A pump is driven by an electric motor running at 1145 rpm and the speed of the compressor is 920rpm. If the diameter of the motor pulley is 160mm, calculate the diameter of the pump pulley


1
Expert's answer
2021-02-15T05:40:17-0500

driving pulley:              driven pulley:

 Diameter = 160mm          Diameter = dmm

rotation = 1145rpm         rotation = 920rpm




(1000×π×160) = 920×π×d

⇒ d = "\frac{1000×π×160}{920×π}"




⇒ d = 173.913mm